# Recursive Neyman Algorithm for Optimum Sample Allocation under Box Constraints on Sample Sizes in Strata

### v3

@@ Abstract
Key Words: Stratified sampling; Optimum sample allocation; Optimum allocation; Optimum allocation under box constraints; Neyman allocation; Recursive Neyman algorithm.
@@ 1 Introduction
[…]
@@ 2 Motivation and literature review
Below, we briefly review existing methods for solving Problem 1.1, including also methods that provide integer-valued solutions.
@@ 2.1 Not-necessarily integer-valued allocation
An approximate solution to Problem 1.1 can be achieved through generic methods of non-linear programming (NLP) (see, e.g. the monograph Valliant et al., 2018, and references therein). These methods have been involved in the problem of optimum sample allocation since solving the allocation problem is equivalent to finding the extreme (namely, stationary points) of a certain objective function over a feasible set. Knowing the (approximate) extreme of the objective function, one can give the (approximate, yet typically sufficiently accurate) sizes of the samples allocated to individual strata.
@@ 2.2 Integer-valued allocation
Integer-valued algorithms dedicated to Problem 1.1 are proposed in Friedrich et al. (2015), Wright (2017, 2020). The multivariate version of the optimum sample allocation problem under box constraints in which $m_{h}=m,\,h\in\mathcal{H}$, for a given constant $m$, is considered in the paper of de Moura Brito et al. (2015). The proposed procedure that solves that problem, uses binary integer programming algorithm and can be applied to the univariate case. See also Brito et al. (2017) for the R-implementation of this approach.
Integer-valued allocation methods proposed in these papers are precise (not approximate) and theoretically sound. However, they are relatively slow, when compared with not-necessarily integer-valued algorithms. For instance, at least for one-sided constraints, the integer-valued algorithm capacity scaling of Friedrich et al. (2015) may be thousands of times slower than the RNA (see Wesołowski et al., 2022, Section 4). This seems to be a major drawback of these methods as the differences in variances of estimators based on integer-rounded non-integer optimum allocation and integer optimum allocation are negligible as explained in Section 6. The computational efficiency is of particular significance when the number of strata is large, see, e.g. application to the German census in Burgard and Münnich (2012), and it becomes even more pronounced in iterative solutions to stratification problems, when the number of iterations may count in millions (see, e.g. Lednicki and Wieczorkowski, 2003, Gunning and Horgan, 2004, Ahmad et al., 2008, Baillargeon and Rivest, 2011, Barcaroli, 2014).
@@ 3 Optimality conditions
$s(\mathcal{L},\,\mathcal{U})=\frac{n-\sum_{h\in\mathcal{L}}m_{h}-\sum_{h\in\mathcal{U}}M_{h}}{\sum_{h\in\mathcal{H}}A_{h}-\sum_{h\in\mathcal{L}\cup\mathcal{U}}A_{h}}.$ (3.1)
@@
[…]
@@
The optimization Problem 1.1 has a unique optimal solution. Point $\mathbf{x}^{*}\in{\mathbb{R}}_{+}^{\lvert\mathcal{H}\rvert}$ is a solution to optimization Problem 1.1 if and only if $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$, with disjoint *$\mathcal{L}^{*},\,\mathcal{U}^{*}\subseteq\mathcal{H}$, such that one of the following two cases holds:
@@ 1. CASE I:
    *$\mathcal{L}^{*}\cup\mathcal{U}^{*}\subsetneq\mathcal{H}$ and
    *=$\begin{gathered}\mathcal{L}^{*}=\left\{h\in\mathcal{H}:\,s(\mathcal{L}^{*},\,\mathcal{U}^{*})\leq\tfrac{m_{h}}{A_{h}}\right\},\\ \mathcal{U}^{*}=\left\{h\in\mathcal{H}:\,s(\mathcal{L}^{*},\,\mathcal{U}^{*})\geq\tfrac{M_{h}}{A_{h}}\right\}.\end{gathered}$ (3.3)
@@ 2. CASE II:
    *$\mathcal{L}^{*}\cup\mathcal{U}^{*}=\mathcal{H}$ and
@@ 4.1 The RNABOX algorithm
In this section we introduce an algorithm solving Problem 1.1. In view of Theorem 3.1 its essential task is to split the set of all strata labels $\mathcal{H}$ into three subsets of take-min ($\mathcal{L}^{*}$), take-max ($\mathcal{U}^{*}$), and take-Neyman ((*$\mathcal{H}\setminus(\mathcal{L}^{*}\cup\mathcal{U}^{*})$). We call this new algorithm RNABOX since it generalizes existing algorithm RNA in the sense that RNABOX solves optimum allocation problem with simultaneous lower and upper bounds, while the RNA is dedicated for the problem with upper bounds only, i.e. for Problem 2.1. Moreover, RNABOX uses RNA in one of its interim steps. We first recall RNA algorithm and then present RNABOX. For more information on RNA, see Wesołowski et al. (2022, Section 2) or Särndal et al. (1992, Remark 12.7.1, p. 466).
@@ 4.2 An example of performance of RNABOX
Table 4.1: An example of RNABOX performance for a population with 10 strata and total sample size $n=5110$. Columns $\mathcal{L}_{r}/\mathcal{U}_{r}$, $r=1,\ldots,6$, represent the content of sets $\mathcal{L},\,\mathcal{U}$ respectively, in the $r$-th iteration of the RNABOX (between Step 4 and Step 5): symbols $\blacksquare$  or $\square$  indicate that a given stratum label $h$ is in $\mathcal{L}_{r}$ or $\mathcal{U}_{r}$, respectively.
$h$
$A_{h}$ $m_{h}$ $M_{h}$ $\mathcal{L}_{1}/\mathcal{U}_{1}$ $\mathcal{L}_{2}/\mathcal{U}_{2}$ $\mathcal{L}_{3}/\mathcal{U}_{3}$ $\mathcal{L}_{4}/\mathcal{U}_{4}$ $\mathcal{L}_{5}/\mathcal{U}_{5}$ $\mathcal{L}_{6}/\mathcal{U}_{6}$ $\mathbf{x}^{*}$
1
2700 750 900 $\square$ $\square$ $\square$ $\square$ $750$
2
2000 450 500 $\blacksquare$ $\square$ $\square$ $\square$ $\square$ $450$
3
4200 250 300 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $261.08$
4
4400 350 400 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\square$ $350$
5
3200 150 200 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $198.92$
6
6000 550 600 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\square$ $\square$ $550$
7
8400 650 700 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\square$ $650$
8
1900 50 100 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $100$
9
5400 850 900 $\blacksquare$ $\blacksquare$ $\square$ $\square$ $\square$ $850$
10
2000 950 1000 $\square$ $\square$ $\square$ $\square$ $\square$ $950$
@@
For this example, RNABOX stops after $6$ iterations with take-min strata set *=$\mathcal{L}^{*}=\{1,2,4,6,7,9,10\}$, take-max strata set $\mathcal{U}^{*}=\{8\}$ and take-Neyman strata set (*$\mathcal{H}\setminus(\mathcal{L}^{*}\cup\mathcal{U}^{*})=\{3,5\}$ (see column $\mathcal{L}_{6}/\mathcal{U}_{6}$). The optimum allocation is a regular one and it is given in column $\mathbf{x}^{*}$ of Table 4.1. The corresponding value of the objective function (1.3) is $441591.5$. The details of interim allocations of strata to sets $\mathcal{L},\,\mathcal{U}$ at each of $6$ iterations of the algorithm are given in columns $\mathcal{L}_{1}/\mathcal{U}_{1}$ $\mathcal{L}_{6}/\mathcal{U}_{6}$.
@@ 4.3.3 Using prior information in RNA at Step 3
$*.\mathcal{U}_{r}=\{h\in\mathcal{H}\setminus\mathcal{L}_{r}:\,x^{**}_{h}=M_{h}\}\subseteq\mathcal{U}_{r-1},\mathrlap{\qquad r=2,\ldots,r^{*}.}$
@@
$*.\mathcal{U}_{r}=\{h\in\mathcal{U}_{r-1}:\,x^{**}_{h}=M_{h}\},\mathrlap{\qquad r=2,\ldots,r^{*}.}$ (4.1)
Given above observation and the fact that from the implementation point of view set $\mathcal{U}_{r}$ is determined internally by RNA, it is tempting to consider modification of RNA such that it makes use of the domain of discourse $\mathcal{U}_{r-1}$ for set $\mathcal{U}_{r}$. This domain could be specified as an additional input parameter, say $\mathcal{J}\subseteq\mathcal{H}$, and then Step 3 of RNA algorithm would read:
@@ 4.4 On a naive extension of the one-sided RNA
1
420 24 420 $30$ $54.44$
2
352 15 88 $88$blacksquare $45.63$
3
2689 1344 2689 $1344$square $1344$square
4
308 8 308 $22$ $39.93$
5
130 3 5 $5$blacksquare $5$blacksquare
SUM
1394 3510 $4891$/2/2 $4891$/1/3
For this population, a naive version of the recursive Neyman stops in the second iteration and it gives allocation $\mathbf{x}$ with =$\mathcal{L}=\{3\}$ and $\mathcal{U}=\{2,5\}$, while the optimum choice is $\mathbf{x}^{*}$ with *=$\mathcal{L}^{*}=\{3\}$ and $\mathcal{U}^{*}=\{5\}$. For non-optimum allocation $\mathbf{x}$, we have $s(\{3\},\,\{2,5\})=0.0714\ngeq 0.25=\tfrac{A_{2}}{M_{2}}$, i.e. $\mathcal{U}$ does not follow (3.3). Note that the allocation $\mathbf{x}$ is feasible in this case (it may not be in general), yet is not a minimizer of the objective function (1.3) as $f(\mathbf{x})=20360-B$, and $f(\mathbf{x}^{*})=17091-B$, where $B$ is some constant. Thus, it appears that in the case of two-sided constraints, one needs a more subtle modification of RNA procedure that lies in a proper update of the pair $(\mathcal{L}_{r},\,\mathcal{U}_{r})$ in each iteration $r=1,\ldots,r^{*}\geq 1$ of the algorithm, such as e.g. in RNABOX.
@@ 5 Numerical results
Figure 5.1: Running times of FPIA and RNABOX for two artificial populations. Top graphs show the empirical median of execution times (calculated from 100 repetitions) for different total sample sizes. Numbers in brackets are the numbers of iterations of a given algorithm. In the case of RNABOX, it is a vector with number of iterations of the RNA (see Step 3 of RNABOX) for each outer iteration of RNABOX. Thus, the length of this vector is equal to the number of outer iterations of RNABOX. Counts of take-min, take-Neyman, and take-max strata are shown on bottom graphs.
From Fig. 5.1 we see that, while for majority of the cases the FPIA is slightly faster than RNABOX, the running times of both of these algorithms are generally comparable. The gain in the execution time of the FPIA results from the fact that it typically runs through a smaller number of sets $\mathcal{L},\,\mathcal{U}\subset\mathcal{H}$ (according to (A.6)), than RNABOX in order to find the optimal $\mathcal{L}^{*}$ and $\mathcal{U}^{*}$. Although this approach usually gives correct results (as in the simulations reported in this section), it may happen that the FPIA misses the optimal sets *$\mathcal{L}^{*},\,\mathcal{U}^{*}\subsetneq\mathcal{H}$. In such a case, FPIA does not converge, as explained in Appendix A.2. Nevertheless, we emphasize that this situation rarely happens in practice.
@@ Appendix A Appendix: On the existing allocation algorithms by GGM and MSW
Table A.1: Two allocations for an example population with two strata: non-optimum $\mathbf{x}^{noptcond}$ with =$\mathcal{L}=\{1\},\,\mathcal{U}=\emptyset$, and optimum $\mathbf{x}^{*}$ with *$\mathcal{L}^{*}=\emptyset,\,\mathcal{U}^{*}=\{1\}$. Set function $s$ is defined in (3.1).
@@ A.2 Fixed-point iteration of MSW
is affine. More specifically, Problem 1.1 is a convex optimization problem of a particular type in which inequality constraint functions (B.2) - (B.3) are affine. See the Appendix E for the definition of the convex optimization problem.
@@
$-\tfrac{A_{h}^{2}}{x^{*2}_{h}}+\lambda-\mu_{h}^{m}+\mu_{h}^{M}=0,\mathrlap{\qquad h\in\mathcal{H},}$ (B.4)
@@
To prove Theorem 3.1, it suffices to show that for $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ with *$\mathcal{L}^{*},\,\mathcal{U}^{*}$ satisfying conditions of CASE I or CASE II, there exist $\lambda\in{\mathbb{R}}$ and $\mu^{m}_{h},\,\mu^{M}_{h}\geq 0,\,h\in\mathcal{H}$, such that (B.4) - (B.8) hold. It should also be noted that the requirement $m_{h}<M_{h},\,h\in\mathcal{H}$, guarantees that $\mathcal{L}^{*}$ and $\mathcal{U}^{*}$ defined in (3.3) and (3.4) are disjoint. Therefore, $\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ is well-defined according to Definition 3.2.

### v4

@@ Abstract
Key Words: Stratified sampling; Optimum allocation; Optimum allocation under box constraints; Neyman allocation; Recursive Neyman algorithm.
@@ 1 Introduction
$x_{h}=\begin{cases}m_{h},&\mathrlap{\qquad h\in\mathcal{L}}\\ M_{h},&\mathrlap{\qquad h\in\mathcal{U},}\end{cases}$
@@ 2 Motivation and literature review
Below, we briefly review existing methods for solving Problem 1.1, including methods that provide integer-valued solutions.
@@ 2.1 Not-necessarily integer-valued allocation
An approximate solution to Problem 1.1 can be achieved through generic methods of non-linear programming (NLP) (see, e.g. the monograph Valliant et al., 2018, and references therein). These methods have been involved in the problem of optimum sample allocation since solving the allocation problem is equivalent to finding the extreme (namely, stationary points) of a certain objective function over a feasible set. Knowing the (approximate) extreme of the objective function, one can determine the (approximate, yet typically sufficiently accurate) sizes of samples allocated to individual strata.
@@ 2.2 Integer-valued allocation
Integer-valued algorithms dedicated to Problem 1.1 are proposed in Friedrich et al. (2015), Wright (2017, 2020). The multivariate version of the optimum sample allocation problem under box constraints in which $m_{h}=m,\,h\in\mathcal{H}$, for a given constant $m$, is considered in the paper of de Moura Brito et al. (2015). The proposed procedure that solves that problem uses binary integer programming algorithm and can be applied to the univariate case. See also Brito et al. (2017) for the R-implementation of this approach.
Integer-valued allocation methods proposed in these papers are precise (not approximate) and theoretically sound. However, they are relatively slow, when compared with not-necessarily integer-valued algorithms. For instance, at least for one-sided constraints, the integer-valued algorithm capacity scaling of Friedrich et al. (2015) may be thousands of times slower than the RNA (see Wesołowski et al., 2022, Section 4). This seems to be a major drawback of these methods as the differences in variances of estimators based on integer-rounded non-integer optimum allocation and integer optimum allocation are negligible as explained in Section 6. The computational efficiency is of particular significance when the number of strata is large, see, e.g. application to the German census in Burgard and Münnich (2012), and it becomes even more pronounced in iterative solutions to stratification problems, when the number of iterations may count in millions (see, e.g. Lednicki and Wieczorkowski, 2003, Gunning and Horgan, 2004, Khan et al., 2008, Baillargeon and Rivest, 2011, Barcaroli, 2014).
@@ 3 Optimality conditions
$s(\mathcal{L},\,\mathcal{U})=\frac{n-\sum_{h\in\mathcal{L}}m_{h}-\sum_{h\in\mathcal{U}}M_{h}}{\sum_{h\in\mathcal{H}\setminus(\mathcal{L}\cup\mathcal{U})}A_{h}}.$ (3.1)
@@
$x_{h}^{(\mathcal{L},\,\mathcal{U})}=\begin{cases}m_{h},&\mathrlap{\qquad h\in\mathcal{L}}\\ M_{h},&\mathrlap{\qquad h\in\mathcal{U}}\\ A_{h}\,s(\mathcal{L},\,\mathcal{U}),&\mathrlap{\qquad h\in\mathcal{H}\setminus(\mathcal{L}\cup\mathcal{U}).}\end{cases}$ (3.2)
@@
The optimization Problem 1.1 has a unique optimal solution. Point $\mathbf{x}^{*}\in{\mathbb{R}}_{+}^{\lvert\mathcal{H}\rvert}$ is a solution to optimization Problem 1.1 if and only if $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$, with disjoint $\mathcal{L}^{*},\,\mathcal{U}^{*}\subseteq\mathcal{H}$, such that one of the following two cases holds:
@@ 1. CASE I:
    $\mathcal{L}^{*}\cup\mathcal{U}^{*}\subsetneq\mathcal{H}$ and
    =$\begin{gathered}\mathcal{L}^{*}=\left\{h\in\mathcal{H}:\,s(\mathcal{L}^{*},\,\mathcal{U}^{*})\leq\tfrac{m_{h}}{A_{h}}\right\},\\ \mathcal{U}^{*}=\left\{h\in\mathcal{H}:\,s(\mathcal{L}^{*},\,\mathcal{U}^{*})\geq\tfrac{M_{h}}{A_{h}}\right\}.\end{gathered}$ (3.3)
@@ 2. CASE II:
    $\mathcal{L}^{*}\cup\mathcal{U}^{*}=\mathcal{H}$ and
@@ 4.1 The RNABOX algorithm
In this section we introduce an algorithm solving Problem 1.1. In view of Theorem 3.1 its essential task is to split the set of all strata labels $\mathcal{H}$ into three subsets of take-min ($\mathcal{L}^{*}$), take-max ($\mathcal{U}^{*}$), and take-Neyman ($\mathcal{H}\setminus(\mathcal{L}^{*}\cup\mathcal{U}^{*})$). We call this new algorithm RNABOX since it generalizes existing algorithm RNA in the sense that RNABOX solves optimum allocation problem with simultaneous lower and upper bounds, while the RNA is dedicated for the problem with upper bounds only, i.e. for Problem 2.1. Moreover, RNABOX uses RNA in one of its interim steps. We first recall RNA algorithm and then present RNABOX. For more information on RNA, see Wesołowski et al. (2022, Section 2) or Särndal et al. (1992, Remark 12.7.1, p. 466).
@@ 4.2 An example of performance of RNABOX
Table 4.1: An example of RNABOX performance for a population with 10 strata and total sample size $n=5110$. Columns $\mathcal{L}_{r}/\mathcal{U}_{r}$, $r=1,\ldots,6$, represent the content of sets $\mathcal{L},\,\mathcal{U}$ respectively, in the $r$-th iteration of the RNABOX (between Step 4 and Step 5): symbols $\square$  or $\blacksquare$  indicate that the stratum with label $h$ is in $\mathcal{L}_{r}$ or $\mathcal{U}_{r}$, respectively.
$h$ $A_{h}$ $m_{h}$ $M_{h}$ $\mathcal{L}_{1}/\mathcal{U}_{1}$ $\mathcal{L}_{2}/\mathcal{U}_{2}$ $\mathcal{L}_{3}/\mathcal{U}_{3}$ $\mathcal{L}_{4}/\mathcal{U}_{4}$ $\mathcal{L}_{5}/\mathcal{U}_{5}$ $\mathcal{L}_{6}/\mathcal{U}_{6}$ $\mathbf{x}^{*}$
1 2700 750 900 $\square$ $\square$ $\square$ $\square$ $750$
2 2000 450 500 $\blacksquare$ $\square$ $\square$ $\square$ $\square$ $450$
3 4200 250 300 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $261.08$
4 4400 350 400 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\square$ $350$
5 3200 150 200 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $198.92$
6 6000 550 600 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\square$ $\square$ $550$
7 8400 650 700 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\square$ $650$
8 1900 50 100 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $100$
9 5400 850 900 $\blacksquare$ $\blacksquare$ $\square$ $\square$ $\square$ $850$
10 2000 950 1000 $\square$ $\square$ $\square$ $\square$ $\square$ $950$
@@
For this example, RNABOX stops after $6$ iterations with take-min strata set =$\mathcal{L}^{*}=\{1,2,4,6,7,9,10\}$, take-max strata set $\mathcal{U}^{*}=\{8\}$ and take-Neyman strata set $\mathcal{H}\setminus(\mathcal{L}^{*}\cup\mathcal{U}^{*})=\{3,5\}$ (see column $\mathcal{L}_{6}/\mathcal{U}_{6}$). The optimum allocation is a regular one and it is given in column $\mathbf{x}^{*}$ of Table 4.1. The corresponding value of the objective function (1.3) is $441591.5$. The details of interim allocations of strata to sets $\mathcal{L},\,\mathcal{U}$ at each of $6$ iterations of the algorithm are given in columns $\mathcal{L}_{1}/\mathcal{U}_{1}$ $\mathcal{L}_{6}/\mathcal{U}_{6}$.
@@ 4.3.3 Using prior information in RNA at Step 3
$\mathcal{U}_{r}=\{h\in\mathcal{H}\setminus\mathcal{L}_{r}:\,x^{**}_{h}=M_{h}\}\subseteq\mathcal{U}_{r-1},\mathrlap{\qquad r=2,\ldots,r^{*}.}$
@@
$\mathcal{U}_{r}=\{h\in\mathcal{U}_{r-1}:\,x^{**}_{h}=M_{h}\},\mathrlap{\qquad r=2,\ldots,r^{*}.}$ (4.1)
Given the above observation and the fact that from the implementation point of view set $\mathcal{U}_{r}$ is determined internally by RNA, it is tempting to consider modification of RNA such that it makes use of the domain of discourse $\mathcal{U}_{r-1}$ for set $\mathcal{U}_{r}$. This domain could be specified as an additional input parameter, say $\mathcal{J}\subseteq\mathcal{H}$, and then Step 3 of RNA algorithm would read:
@@ 4.4 On a naive extension of the one-sided RNA
1 420 24 420 $30$ $54.44$
2 352 15 88 $88$blacksquare $45.63$
3 2689 1344 2689 $1344$square $1344$square
4 308 8 308 $22$ $39.93$
5 130 3 5 $5$blacksquare $5$blacksquare
SUM 1394 3510 $4891$/2/2 $4891$/1/3
For this population, a naive version of the recursive Neyman stops in the second iteration and it gives allocation $\mathbf{x}$ with =$\mathcal{L}=\{3\}$ and $\mathcal{U}=\{2,5\}$, while the optimum choice is $\mathbf{x}^{*}$ with =$\mathcal{L}^{*}=\{3\}$ and $\mathcal{U}^{*}=\{5\}$. For non-optimum allocation $\mathbf{x}$, we have $s(\{3\},\,\{2,5\})=0.0714\ngeq 0.25=\tfrac{A_{2}}{M_{2}}$, i.e. $\mathcal{U}$ does not follow (3.3). Note that the allocation $\mathbf{x}$ is feasible in this case (it may not be in general), yet is not a minimizer of the objective function (1.3) as $f(\mathbf{x})=20360-B$, and $f(\mathbf{x}^{*})=17091-B$, where $B$ is some constant. Thus, it appears that in the case of two-sided constraints, one needs a more subtle modification of RNA procedure that lies in a proper update of the pair $(\mathcal{L}_{r},\,\mathcal{U}_{r})$ in each iteration $r=1,\ldots,r^{*}\geq 1$ of the algorithm, such as e.g. in RNABOX.
@@ 5 Numerical results
Figure 5.1: Running times of FPIA and RNABOX for two artificial populations. Top graphs show the empirical median of execution times (calculated from 100 repetitions) for different total sample sizes. Numbers in brackets are the numbers of iterations of a given algorithm. In the case of RNABOX, it is a vector with number of iterations of the RNA (see Step 3 of RNABOX) for each iteration of RNABOX. Thus, the length of this vector is equal to the number of iterations of RNABOX. Counts of take-min, take-Neyman, and take-max strata are shown on bottom graphs.
From Fig. 5.1 we see that, while for majority of the cases the FPIA is slightly faster than RNABOX, the running times of both of these algorithms are generally comparable. The gain in the execution time of the FPIA results from the fact that it typically runs through a smaller number of sets $\mathcal{L},\,\mathcal{U}\subset\mathcal{H}$ (according to (A.6)), than RNABOX in order to find the optimal $\mathcal{L}^{*}$ and $\mathcal{U}^{*}$. Although this approach usually gives correct results (as in the simulations reported in this section), it may happen that the FPIA misses the optimal sets $\mathcal{L}^{*},\,\mathcal{U}^{*}\subsetneq\mathcal{H}$. In such a case, FPIA does not converge, as explained in Appendix A.2. Nevertheless, we emphasize that this situation rarely happens in practice.
@@ Appendix A Appendix: On the existing allocation algorithms by GGM and MSW
Table A.1: Two allocations for an example population with two strata: non-optimum $\mathbf{x}^{noptcond}$ with =$\mathcal{L}=\{1\},\,\mathcal{U}=\emptyset$, and optimum $\mathbf{x}^{*}$ with $\mathcal{L}^{*}=\emptyset,\,\mathcal{U}^{*}=\{1\}$. Set function $s$ is defined in (3.1).
@@ A.2 Fixed-point iteration of MSW
is affine. More specifically, Problem 1.1 is a convex optimization problem of a particular type in which inequality constraint functions (B.2) - (B.3) are affine. See Appendix E for the definition of the convex optimization problem.
@@
As explained in Remark B.1, Problem 1.1 is a convex optimization problem in which the inequality constraint functions $g_{h}^{m},\,g_{h}^{M},\,h\in\mathcal{H}$ are affine. The optimal solution for such a problem can be identified through the Karush-Kuhn-Tucker (KKT) conditions, in which case they are not only necessary but also sufficient; for further references, see Appendix E.
The gradients of the objective function (B.1) and constraint functions (B.2) - (B.3) are as follows:
@@
$-\tfrac{A_{h}^{2}}{{(x^{*}_{h})}^{2}}+\lambda-\mu_{h}^{m}+\mu_{h}^{M}=0,\mathrlap{\qquad h\in\mathcal{H},}$ (B.4)
@@
To prove Theorem 3.1, it suffices to show that for $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ with $\mathcal{L}^{*},\,\mathcal{U}^{*}$ satisfying conditions of CASE I or CASE II, there exist $\lambda\in{\mathbb{R}}$ and $\mu^{m}_{h},\,\mu^{M}_{h}\geq 0,\,h\in\mathcal{H}$, such that (B.4) - (B.8) hold. It should also be noted that the requirement $m_{h}<M_{h},\,h\in\mathcal{H}$, guarantees that $\mathcal{L}^{*}$ and $\mathcal{U}^{*}$ defined in (3.3) and (3.4) are disjoint. Therefore, $\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ is well-defined according to Definition 3.2.